OpenStudy (anonymous):

How to solve X^2+10x=24 by completing the square.

OpenStudy (anonymous):

what times what equals 24 and those same numbers have to add upto 10

OpenStudy (anonymous):

12 and -2 right?

OpenStudy (anonymous):

or 4 and 6

OpenStudy (anonymous):

\[x^2+10x=(x+5)^2-25\] Do you understand this?

OpenStudy (anonymous):

no 4 and 6 when u multiply its 24 not -24

OpenStudy (anonymous):

srry it has to multiply to -24

OpenStudy (anonymous):

... or we could actually help him with the mechanics of the question

OpenStudy (anonymous):

*or her

OpenStudy (anonymous):

so the answer is (x-2)(x+10)

OpenStudy (anonymous):

if you multiply using foils u can see it equals the question

OpenStudy (anonymous):

Really, that's not completing the square.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Completing_the_square Needs to be in the form (x+k)^2=c

OpenStudy (anonymous):

okie

OpenStudy (anonymous):

ur right srry

OpenStudy (anonymous):

Okay, @lizzy87 , do you understand why\[x^2+10x=(x+5)^2-25\]?

OpenStudy (anonymous):

how didi you get the 5?

OpenStudy (anonymous):

 Because 2x5=10